# Spatially covariant gravity with two degrees of freedom in the presence of an auxiliary scalar field: Hamiltonian analysis

### v1

@@ V.2 The first condition
$A\left[\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{1}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 , (96)
$A\left[\frac{2}{N}\frac{\partial c_{1}}{\partial N}+\frac{\partial^{2}c_{1}}{\partial N^{2}}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial N}-\alpha^{2}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{2}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 ,
$A\left[\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{1}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0,0 ,
$A\left[\frac{2}{N}\frac{\partial c_{3}}{\partial N}+\frac{\partial^{2}c_{3}}{\partial N^{2}}-\alpha^{2}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{2}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0.0 . (97)
As long as equations (95) and (96)-(97) are satisfied, the coefficient of $B^{2}$ will be automatically zero. Therefore, we do not have to discuss the differential equation corresponding to the coefficient of $B^{2}$ here.
@@
where the coefficients $C$ and $D$ are both constants.
By plugging the expressions of $c_{1}$ and $c_{3}$ into (96)-(97), we can simplify the four differential equations as
$\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (101)
$\alpha^{2}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (102)
$\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0,0 , (103)
$\alpha^{2}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0.0 . (104)
@@
As we can see, if both $c_{1}$ and $c_{3}$ are independent of $\phi$, constraint $\mathcal{C}_{2}$ will not exist. The number of second-class constraints will be reduced by one, and the number of degrees of freedom will increase by 1/2. This is the linear-constraints branch of the 2DOF condition mentioned in the previous section, which is not the case we are interested in. Therefore, we assume that $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously. As a result, the coefficient $\alpha^{2}$ must be fixed to zero.
@@ V.3 The second condition
According to the above discussion, after imposing the first 2DOF condition $\mathcal{S}_{1}$, we can obtain the constraint $\mathcal{C}^{\prime}$ from the null eigenvector $\mathcal{V}_{a}(\vec{x})$,
@@
The Poisson brackets between the constraint $\mathcal{C}^{\prime}$ and the constraints $p$, $\pi$, and $p^{kl}$ are
@@
which all vanish on the constraint surface. The Poisson bracket between the constraint $\mathcal{C}^{\prime}$ and itself is generally nonzero, unless
@@
As mentioned above, $c_{1}$ and $c_{3}$ cannot both be independent of $\phi$ simultaneously; otherwise, the number of degrees of freedom will increase by 1/21/21 / 2. If we require the second 2DOF condition $\mathcal{S}_{2}^{(2)}$ to hold, there are two cases.
In the first case, if we choose $\frac{\partial c_{1}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
@@
This case is clearly non-physical, as it fixes $R$ to vanish.
On the other hand, if we choose $\frac{\partial c_{3}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
@@
In this case, we can solve for all the coefficients:
@@
The resulting theory thus corresponds to a SCG with an auxiliary scalar field of $d=2$ that propagates two degrees of freedom. The constraint coefficients $\alpha^{i}$ in the previous differential equation are also uniquely determined:
@@
whose determinant is obviously zero. Based on the discussion of null eigenvectors in Appendix C, we can find a null eigenvector $\mathcal{V}_{a}^{(\mathcal{E})}$ to construct a new first-class constraint $\tilde{\mathcal{C}}$,
@@
#$\#_{\text{dof}}$ $=\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (118)
@@
=2=2.= 2 .
Finally, let us discuss another 2DOF condition $\mathcal{S}_{2}^{(1)}$. By applying the consistency condition of the constraint $\mathcal{C}_{2}$, we can obtain the constraint $\Phi$,
@@
Therefore, requiring the 2DOF condition $\mathcal{S}_{2}^{(1)}=\det\mathcal{D}^{ab}\approx 0$ also implies that $\frac{\partial c_{3}}{\partial\phi}$ may vanish as one of the conditions for the matrix $\mathcal{D}^{ab}$ to be degenerate, which is sufficient to satisfy the 2DOF condition $\mathcal{S}_{2}^{(2)}$.
@@ VI Conclusion
In this work, we investigated a class of spatially covariant gravity theories with a non-dynamical scalar field. We briefly described our model in Sec. II. The general Lagrangian is given by (3), which has been shown to have two tensorial and one scalar DOFs [85]. The purpose of this work is to determine the conditions on the Lagrangian, which we dub the 2DOF conditions, under which only two degrees of freedom propagate.
@@
A perturbative approach was taken in [86] to derive conditions such that the scalar mode is eliminated at linear order in perturbations. In this work, we employ a strict Hamiltonian constraint analysis to derive conditions such that only two DOFs propagate in the nonperturbative sense. In Sec. III, we describe the Hamiltonian formalism for the primary constraints and their consistency conditions. As expected, there are 3 DOFs in the theory if no further conditions are imposed.
@@
In Sec. IV, by requiring the degeneracy of the Dirac matrix, we find that two conditions are required to make the theory propagate only 2 DOFs. The first condition (38) states that the matrix $\mathcal{B}^{ab}$ given in (39) must be degenerate, which results in a secondary constraint. To fully eliminate the scalar mode, a second condition is necessary. The second condition can be divided into two categories according to their different effects on the Dirac matrix. In the first case, the condition (49) generally turns a second-class constraint into a first-class constraint. In the second case, the condition (50) generates another secondary constraint $\Phi$.
@@
We would like to emphasize that even if the Lagrangian satisfies both 2DOF conditions, the number of DOFs does not necessarily decrease. A special case is when the constraints are linearly dependent, which implies that the number of DOFs may remain unchanged. Another special case, which we refer to as the non-physical branch of the 2DOF conditions, renders the theory trivial or physically unacceptable (e.g., lacking tensor modes). One must therefore be careful when dealing with the 2DOF conditions in order to select the physical case, i.e., theories with precisely two degrees of freedom.
@@
In Sec. V, based on the spatially covariant monomials classified in [57], [58], we consider a concrete Lagrangian (60) built from monomials of $d=2$ (i.e., with two derivatives) as an illustration of our formal analysis. Since the Dirac matrix of this model contains derivatives of the delta function, it is complicated to calculate the matrix form of the 2DOF conditions for this concrete model. Instead, we choose an equivalent but more efficient method to obtain the 2DOF conditions, which is to find the null eigenvector $\mathcal{V}_{a}$ satisfying (40). The resulting 2DOF conditions are given in (66) and (111). In particular, for this concrete model, we are able to eliminate the non-physical branch and the linear constraints branch of the 2DOF conditions. Finally, we find a set of coefficients (114), and the resulting Lagrangian corresponds to a spatially covariant gravity theory with an auxiliary scalar field that propagates only two DOFs.
@@ Appendix A Comparison with scalar-tensor theory in the spatial gauge
The spatially covariant gravity with an auxiliary scalar field was firstly proposed in [85], where the Hamiltonian analysis was performed to show that the theory propagates 3 degrees of freedom. This idea was originally motivated by generally covariant scalar-tensor theory when the scalar field possesses a spacelike gradient. After choosing the so-called “spatial gauge” as in [85], the resulting action takes a form similar to (1). However, we should note that “spatially covariant gravity with an auxiliary scalar field” and “scalar-tensor theory in the spatial gauge” are completely different theories.
First, we will show the difference between (1) and the scalar-tensor theory in the spatial gauge. Let us consider a general action of scalar-tensor theory
@@
where the Lagrangian involves a scalar field $\phi$, the spacetime metric $g_{ab}$, the spacetime curvature tensor ${}^{4}\!R_{abcd}$ as well as their covariant derivatives. The 4-dimensional Levi-Civita tensor $\varepsilon_{abcd}$ encodes possible parity violation effects.
In a cosmological context, the scalar field $\phi$ is assumed to possess a timelike gradient so that the so-called unitary gauge with $\phi=t$ can be chosen. Here we consider the opposite situation by assuming that the scalar field possesses a spacelike gradient [^3] Similar ideas appear in “elastic inflation” [87] and “solid inflation” [88], where it is the Nambu-Goldstone boson breaking spatial diffeomorphism that plays the role of the scalar field.. Contrary to the unitary gauge which is defined by requiring $\mathrm{D}_{a}\phi=0$, we now choose a gauge in which
£$\pounds_{\bm{n}}\phi=0,$ (122)
where $n^{a}$ is the normal vector to the hypersurfaces. This can be understood as choosing hypersurfaces such that the normal vector $n^{a}$ exactly lies on the constant $\phi$ hypersurfaces. As a result, the value of $\phi$ does not change when transported along the normal vector. This choice of spatial hypersurfaces is referred to as the “spatial gauge” in [85].
The normal vector $n^{a}$ is normalized by $n^{a}n_{a}=-1$. The induced metric on the spatial hypersurfaces is, as usual,
@@
where $\mathrm{D}_{a}$ is the covariant derivative compatible with the induced metric $h_{ab}$. In the last equality of (124) we have used the fact that £$\pounds_{\bm{n}}\phi=0$ in the spatial gauge. This is exactly the opposite of the unitary gauge, where $\mathrm{D}_{a}\phi=0$ and thus $\nabla_{a}\phi\rightarrow-n_{a}\pounds_{\bm{n}}\phi$.
After performing the 3+1 decomposition and choosing the spatial gauge, the action (121) can be written in the form
@@
where “s.g.” stands for the spatial gauge and we have fixed the spatial coordinates adapted to the spacelike hypersurfaces. At first glance, the action (125) takes the same form as (1); however, the crucial difference is that in the spatial gauge, the scalar field $\phi$ can be viewed as a time-independent but space-dependent field, which breaks spatial diffeomorphism. In other words, $\phi=\phi(\bm{x})$ in (125) cannot be viewed as a dynamical nor auxiliary variable, but is actually a “function” of spatial coordinates with fixed values. This is completely different from (1), which has spatial covariance and where $\phi$ plays the role of an auxiliary field.
Next, we show that the generally covariant version (i.e., the correspondence) of (1) is nothing but a “two-field” scalar-tensor theory. Let us take $\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi$ as an example. The covariant version is
@@
where $u_{a}=-N\nabla_{a}\Phi$ with $\Phi$ being the scalar field defining the spacelike hypersurfaces. In other words, the spacelike hypersurfaces are defined by $\Phi=\mathrm{const.}$. In (126), $N=1/\sqrt{-\left(\nabla\Phi\right)^{2}}$, which reduces to the lapse function when fixing the so-called unitary gauge with $\Phi=t$. By expanding (126) explicitly, we get
$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi\rightarrow\left(g^{ab}-\frac{\nabla^{a}\Phi\nabla^{b}\Phi}{\left(\nabla\Phi\right)^{2}}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (127)

### v2

@@ V.2 The first condition
$A\left[\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{1}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big(\frac{\partial c_{1}}{\partial\phi}\Big)^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ $0,$ (96)
$A\left[\frac{2}{N}\frac{\partial c_{1}}{\partial N}+\frac{\partial^{2}c_{1}}{\partial N^{2}}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial N}-\alpha^{2}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{2}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ $0,$
$A\left[\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{1}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ $0,$
$A\left[\frac{2}{N}\frac{\partial c_{3}}{\partial N}+\frac{\partial^{2}c_{3}}{\partial N^{2}}-\alpha^{2}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{2}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ $0.$ (97)
As long as equations (95) and (96)-(97) are satisfied, the coefficient of $B^{2}$ will be automatically fixed to be zero. Therefore, we do not have to discuss the differential equation corresponding to the coefficient of $B^{2}$ here.
@@
where the coefficients $C$ and $D$ are both constant.
By plugging the expressions of $c_{1}$ and $c_{3}$ in (96)-(97), we can simplify the four differential equations
$\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big(\frac{\partial c_{1}}{\partial\phi}\Big)^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}$ $=$ $0,$ (101)
$\alpha^{2}\frac{\partial c_{1}}{\partial\phi}$ $=$ $0,$ (102)
$\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}$ $=$ $0,$ (103)
$\alpha^{2}\frac{\partial c_{3}}{\partial\phi}$ $=$ $0.$ (104)
@@
As we can see, if both $c_{1}$ and $c_{3}$ are independent of $\phi$, constraint $\mathcal{C}_{2}$ will do not exist. The number of second-class constraints will be reduced by one, and the number of degrees of freedom will be increased by 1/2. This is the linear constraints branch of 2DOF condition we mentioned in the previous section, which is not the case we are interested in. Therefore we assume that $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously. As a result, the coefficient $\alpha^{2}$ has to be fixed to be zero.
@@ V.3 The second condition
According to the above discussion, after requiring the first 2DOF condition $\mathcal{S}_{1}$, we can obtain constraint $\mathcal{C}^{\prime}$ by the null eigenvector $\mathcal{V}_{a}(\vec{x})$,
@@
Poisson brackets between the constraint $\mathcal{C}^{\prime}$ and constraints $p$, $\pi$, $p^{kl}$ are
@@
which all vanish on the constraint surface. Poisson bracket between constraint $\mathcal{C}^{\prime}$ and itself is generally nonzero, unless
@@
As we have mentioned above, $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously, otherwise the number of degrees of freedom will increased by $1/2$. If we require the second 2DOF condition $\mathcal{S}_{2}^{(2)}$ to be valid, there will be two cases.
In the first case, if we choose $\frac{\partial c_{1}}{\partial\phi}=0$, constraint $\mathcal{C}_{2}$ will become
@@
This case is clearly non-physical as it fixes $R$ to be vanishing.
On the other hand, if we choose $\frac{\partial c_{3}}{\partial\phi}=0$, constraint $\mathcal{C}_{2}$ will become
@@
In this case, we can solve all the coefficients
@@
The resulting theory thus corresponds to a SCG with an auxiliary scalar field of $d=2$ that propagates two degrees of freedom. The constraint coefficients $\alpha^{i}$ in the previous differential equation are also uniquely determined
@@
whose determinant is obviously zero. Based on the discussion on the null eigenvectors in Appendix (C), we can find a null eigenvector $\mathcal{V}_{a}^{(\mathcal{E})}$ to build a new first-class constraint $\tilde{\mathcal{C}}$,
@@
$\#_{\text{dof }}$ $=\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (118)
@@
$=2.$
Finally let us discuss another 2DOF condition $\mathcal{S}_{2}^{(1)}$. By applying the consistency condition of constraint $\mathcal{C}_{2}$, we can obtain the constraint $\Phi$,
@@
So when we require the 2DOF condition $\mathcal{S}_{2}^{(1)}=\det\mathcal{D}^{ab}\approx 0$, it will also imply that $\frac{\partial c_{3}}{\partial\phi}$ may be vanishing as one of the conditions for the matrix $\mathcal{D}^{ab}$ to be degenerate, which is sufficient to satisfy 2DOF condition $\mathcal{S}_{2}^{(2)}$.
@@ VI Conclusion
In this work, we investigated a class of spatially covariant gravity theories with a non-dynamical scalar field. We briefly describe our model in Sec. II. The general Lagrangian is given by (3), which has been proved to have two tensorial and one scalar DOFs [85]. The purpose of this work is to determine the conditions on the Lagrangian, which we dub the 2DOF conditions, under which only two degrees of freedom are propagating.
@@
A perturbative approach has been taken in [86] to derive conditions such that the scalar mode is eliminated at linear order in perturbations. In this work we employ the strict Hamiltonian constraint analysis to derive the conditions such that that only two DOFs are propagating in the nonperturbative sense. In Sec. III, we describe the Hamiltonian formalism for the primary constraints and their consistency conditions. As expected, there are 3 DOFs in the theory if no further conditions are imposed.
@@
In Sec. IV, by requiring the degeneracy of the Dirac matrix, we find that two conditions are required to make the theory propagate only 2DOFs. The first condition (38) states that the matrix $\mathcal{B}^{ab}$ given in (39) must be degenerate, which will result in a secondary constraint. In order to fully eliminate the scalar mode, a second condition is necessary. The second condition can be divided into two categories according to their different effects on the Dirac matrix. In the first case, the condition (49) will generally turn a second-class constraint into a first-class constraint. In the second case, the condition (50) will generate another secondary constraint $\Phi$.
@@
We would like to emphasize that even the Lagrangian satisfies the both two 2DOF conditions, the number of DOFs do not necessarily decrease. A special case is that the constraints are linearly dependent, which implies that the number of DOFs may remain unchanged. Another special case, which we refer to as the non-physical branch of the 2DOF conditions, will make the theory trivial or physically unacceptable (e.g., without tensor modes). One thus must be careful when dealing with the 2DOF conditions in order to pick up the physical case, i.e., theories with precisely two degrees of freedom.
@@
In Sec. V, based on the spatially covariant monomials classified in [57, 58], we consider a concrete Lagrangian (60) built of monomials of $d=2$ (i.e., with two derivatives) as an illustration of our formal analysis. Since the Dirac matrix of this model contains derivatives of the delta function, it is complicated to calculate the matrix form of the 2DOF conditions of this concrete model. Instead, we choose an equivalent but more efficient way to obtain the 2DOF conditions, which is to find out the null eigenvector $\mathcal{V}_{a}$ satisfing (40). The resulting 2DOF conditions are given in (66) and (111). In particular, for this concrete model, we are able to eliminate the non-physical branch and the linear constraints branch of the 2DOF conditions. Finally, we find a set of coefficients (114), and the resulting Lagrangian corresponds to a spatially covariant gravity theory with an auxiliary scalar field which propagates only two DOFs.
@@ Appendix A Comparison with scalar-tensor theory in the spatial gauge
The spatially covariant gravity the an auxiliary scalar field was firstly proposed in [85], in which the Hamiltonian analysis was performed in order to show that the theory propagates 3 degrees of freedom. This idea was originally motivated by generally covariant scalar-tensor theory when the scalar field possesses a spacelike gradient. After choosing the so-called “spatial gauge” as in [85], the resulting action takes the form that is similar to (1). However, we should notice that “spatially covariant gravity the an auxiliary scalar field” and “scalar-tensor theory in the spatial gauge” are completely different theories.
First we will show the difference between (1) and the scalar-tensor theory in the spatial gauge. Let us consider a general action of scalar-tensor theory
@@
where the Lagrangian involves a scalar field $\phi$, spacetime metric $g_{ab}$, the spacetime curvature tensor $^{4}\!R_{abcd}$ as well as their covariant derivatives. The 4-dimension Levi-Civita tensor $\varepsilon_{abcd}$ encodes possible parity violation effects.
In cosmological context, the scalar field $\phi$ is assumed to possesses a timelike gradient so that the so-called unitary gauge with $\phi=t$ can be chosen. Here we consider the contrary situation by assuming that the scalar field possesses a spacelike gradient [^3] Similar ideas appear in “elastic inflation” [87] and “solid inflation” [88] where it is the Nambu-Goldstone boson breaking spatial diffeomorphism that plays the role of the scalar field.. Contrary to the unitary gauge which is defined by requiring $\mathrm{D}_{a}\phi=0$, we now can choose a gauge in which
$\pounds_{\bm{n}}\phi=0,$ (122)
where $n^{a}$ is the normal vector to the hypersurfaces. This can be understood that we choose hypersurfaces such that the normal vector $n^{a}$ exactly lie on the the constant $\phi$ hypersurfaces. As a result, the value of $\phi$ does not change when being transported along the normal vector. This choice of spatial hypersurfaces is referred as “spatial gauge” in [85].
The normal vector $n^{a}$ is normalized by $n^{a}n_{a}=-1$. The induced metric on the spatial hypersurfaces is as usual
@@
where $\mathrm{D}_{a}$ is the covariant derivative compatible with induced metric $h_{ab}$. In the last equality of (124) we have used the fact that $\pounds_{\bm{n}}\phi=0$ in the spatial gauge. This is exactly contrary to what in the unitary gauge where $\mathrm{D}_{a}\phi=0$ and thus $\nabla_{a}\phi\rightarrow-n_{a}\pounds_{\bm{n}}\phi$.
After making the 3+1 decomposition and choosing the spatial gauge, the action (121) can written in the form
@@
where “s.g.” stands for the spatial gauge and we have fixed the spatial coordinates adapted to the spacelike hypersurfaces. At the first glance, the action (125) takes the same form as (1), however, the crucial difference is that in the spatial gauge, the scalar field $\phi$ can be viewed as a time-independent but space-dependent field, which breaks spatial diffeomorphism. In other words, $\phi=\phi(\bm{x})$ in (125) cannot be viewed as a dynamical nor auxiliary variable, which is actually a “function” of space coordinates with fixed values. This is completely different from (1) which has spatial covariance, in which $\phi$ plays the role of an auxiliary field.
Then we show that the generally covariant version (i.e., correspondence) of (1) is nothing but a “two-field” scalar-tensor theory. Let us take $\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi$ as an example. The covariant version is
@@
where $u_{a}=-N\nabla_{a}\Phi$ with $\Phi$ is the scalar field defining the spacelike hypersurfaces. In other words, the spacelike hypersurfaces are defined by hypersurfaces with $\Phi=\mathrm{const.}$. In (126), $N=1/\sqrt{-\left(\nabla\Phi\right)^{2}}$, which reduces to the lapse function when fixing the so-called unitary gauge with $\Phi=t$. By expanding (126) explicitly, we get
$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi\rightarrow\left(g^{ab}-\frac{\nabla^{a}\Phi N\nabla^{b}\Phi}{\left(\nabla\Phi\right)^{2}}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (127)